Delta-Rings and Axioms for Witt Vectors with Frobenius

Luc Bélair

We discuss the role of the delta-ring structure in the model theory of Witt vectors with the Frobenius automorphism.

Address: UQAM, C.P. 8888, Succursale Centre-ville Montréal, Québec, H3C 3P8

E-mail: belair.luc@uqam.ca